# The impact of Solar magnetic field configurations on the production of gamma rays at the Solar disk

### v1

@@ Abstract
The Sun produces a steady signal of high-energy gamma rays through interactions of Galactic cosmic rays (GCRs) with its atmosphere. Observations with Fermi-LAT and HAWC have revealed a gamma-ray flux significantly higher than early theoretical predictions, with unexpected temporal and spectral features that suggest a crucial role of the solar magnetic field. In this work, we model GCR-induced gamma-ray emission at the solar disk using the CRPropa framework with realistic hadronic interactions, chromospheric density profiles, and several magnetic field configurations over the solar cycle. This allows us to quantify the gamma-ray emission of the entire solar disk for different phases of the solar activity cycle and we present, for the first time, maps of the production locations of gamma rays on the solar surface. We consider both mono-energetic and realistic power-law injection spectra in a simplified dipole-quadrupole current sheet model and potential-field source surface (PFSS) extrapolations for Carrington rotations during solar maximum and minimum. Our results show that magnetic mirroring and large-scale field topology strongly affect the spectral shape and spatial distribution of the emission, with slightly enhanced fluxes predicted at solar minimum. While our simulated baseline fluxes remain below observations, additional effects, such as heavier nuclei, Parker-field mirroring, and deeper atmospheric interactions, could result in further enhancements of fluxes closer to observational values. Hadronic interactions do not only produce gamma rays but also neutrinos. We estimate the expected neutrino flux from the Sun based on our gamma-ray predictions. We find that the expected flux is slightly below current upper limits from IceCube.
@@ 1 Introduction
The dominant high-energy cosmic ray component in the heliosphere is of Galactic origin. These Galactic Cosmic Rays (GCRs) are not only useful messengers from distant astrophysical systems (e.g., Becker-Tjus-Merten-2020), but they can also be used to infer information about the large-scale structure of the heliosphere (e.g., Pogorelov-etal-2017) as well as about the turbulence in the solar wind (e.g., Oughton-Engelbrecht-2021; Fraternale-etal-2022).
@@
Even in the innermost heliosphere, high-energetic GCRs are important agents regarding the solar magnetic field. The deflection of GCRs in the latter produces the so-called solar cosmic-ray shadow (Becker-Tjus-etal-2020; Aartsen-etal-2021, and references therein), which varies with solar activity and can, therefore, be used to infer information about the solar coronal magnetic field (e.g., Alfaro-etal-2024; Ng-etal-2025). The GCRs also interact with the atmosphere and the interior of the Sun. Seckel-etal-1991 were the first to estimate the fluxes of gamma-rays, neutrinos, antiprotons, neutrons, and antineutrons resulting from cascades that are initiated by the interactions of GCRs with the solar atmosphere. Aside from protons and heavier ions, GCR electrons and positrons can also produce gamma-rays via inverse Compton scattering of solar photons, as well as extreme ultraviolet (EUV) to hard X-ray emission via synchrotron radiation in the solar magnetic field. The transport of GCR electrons from 1 AU to the Sun and their subsequent synchrotron emission were recently modeled (Petrosian-etal-2023; Orlando-etal-2023). The steady flux of gamma-rays produced near the Sun as a result of GCR proton interactions is the primary interest in the present study.
@@
Gamma rays produced by the Sun in the GeV range are also associated with transient solar flare events. The maximum energy of flare-produced gamma rays is about 4 GeV (e.g. Ackermann-etal-2014), which is lower than most of the energy range discussed in the following, reaching well into the TeV regime. Regarding the steady gamma-ray flux of the Sun, one distinguishes two components of the GCR-induced gamma-ray emission, namely the disk and the halo component. The first one comes from the solar disk and was theoretically studied by Seckel-etal-1991 and detected first in EGRET data (Orlando-Strong-2008). It results predominantly from GCR protons, which follow magnetic flux tubes in the solar atmosphere. When these particles interact with the solar atmosphere, they produce, via cascades, gamma-rays from the direction of the solar disk. The second, so-called halo component, results from inverse Compton scattering of cosmic-ray electrons in the extended solar atmosphere, quantified and validated by Orlando-Strong-2008 as far as possible with the EGRET data available at that time (see also Moskalenko-etal-2006). The predictions for the halo component and the corresponding EGRET observations have been confirmed with measurements made with the Large Area Telescope of the Fermi mission (Fermi-LAT, Abdo-etal-2011). The disk component was detected with Fermi as well, however, at a flux level about a factor of six higher (Nisa-etal-2019) than predicted by Seckel-etal-1991. These findings have been confirmed at lower energies in Ng-etal-2016 by extending the analysis of Fermi-LAT data from 1–10 GeV (Abdo-etal-2011) to the interval 1–100 GeV. With the High Altitude Water Cherenkov (HAWC) Observatory (DeYoung-2012), solar disk gamma rays with higher energies in a low TeV range have been observed (Albert-etal-2023).
@@
Ng-etal-2016 identified a time-dependence of the flux of the disk component by a factor of two to three in the energy interval 1101-10 GeV, confirmed by Linden-etal-2022. This variation is present up to about 30 MeV, anti-correlated to solar activity, and cannot be explained exclusively with the corresponding modulation of GCRs, which amounts to $\sim$15% for 10 GeV protons and less for higher energies. So, as for the cosmic-ray shadow, the solar magnetic field must be expected to play a significant role: The field structure determines the trajectories of the GCRs in the solar atmosphere. First, it determines which particles initiate cascades that eventually lead to the formation of the disk component of gamma rays. And, second, the field also determines the temporal variation of the gamma-ray signal. Such variation with solar activity is also present in the HAWC data (Albert-etal-2023).
@@
Furthermore, Tang-etal-2018 reported an unexpected dip in the solar gamma-ray spectrum between 305030-50 GeV, which has recently been attributed to an energy cutoff in the turbulence-induced migration of GCRs into closed magnetic arcades (Puzzoni-etal-2024). While the confirmation of this needs further measurements, another finding remains, so far, unexplained. Linden-etal-2018 reported that the gamma-ray emission is not point-like and isotropic (Nisa-etal-2019) but exhibits a nearly constant polar component up to 100 GeV and an equatorial component that is brightest during solar minimum with energies above 200 GeV but not in the TeV-regime (Albert-etal-2018; Bartoli-etal-2019). Finally, the gamma-ray flux is characterized by power laws in energy, namely $E^{-2.2}$ below 200 GeV (Linden-etal-2022, and references therein), steepening to $E^{-3.6}$ above 500 GeV (Albert-etal-2023).
@@
Following the original idea by Seckel-etal-1991, recently Li-etal-2024 have attempted to identify the magnetic field structures at the solar surface that are crucial for reflecting hadronic GCRs and have modeled the corresponding gamma-ray emission. To this end, the authors considered a solar surface magnetic field consisting of vertical flux tubes, representing those that form the so-called network elements, and of vertical flux sheets, representing those at the so-called down-flow lanes between granules. This way, the discrepancy between simulations and observations could be somewhat reduced, but it remains at a factor of two to four. While this model is quite advanced on a small scale, it assumes a spherically symmetric solar atmosphere on a large scale and, thus, does not take into account the large-scale coronal and interplanetary magnetic field structure. In the context of gamma-ray production, the coronal field very close to the Sun was taken into account by Hudson-MacKinnon-2020 using an MHD model and by Li-Ng-etal-2024 employing the potential field source surface (PFSS) approach, again resulting in too low simulated fluxes.
@@
All of the above models did not consider a small-scale turbulent component of the large-scale solar magnetic field. This was done recently by Puzzoni-etal-2024, focusing on closed turbulent magnetic field arcades in the low photosphere on the gamma ray flux. This study was complemented by Puzzoni-etal-2025 with modeling a synthetic large-scale magnetic field with static, laminar open field lines in the chromosphere that are increasingly braided near the solar surface. A key finding of these studies is that the tension between the modelled and the observed gamma ray flux from the solar disk reduces with increasing turbulence level. In extension of their earlier work, Li-etal-2025 also included the effect of turbulence. Interestingly, the Alfvénic turbulence considered therein was found to reduce rather than increase the modelled flux.
@@
Extending the scale to a few solar radii, Hutchinson-etal-2022 considered the interplanetary Parker field and the turbulence-induced scattering of GCRs within, however, only in the context of solar flares and not yet computing gamma ray fluxes.
@@
2 Modelling Framework
We model the gamma-ray emission with CRPropa 3.2 (CRPropa3.2), a publicly available tool for the propagation and interaction of charged particles in magnetic fields. While the code was originally written for ultra high energy cosmic rays (UHECRs) (CRPropa3), an extension for lower energies to cover GCRs has been introduced later on (CRPropa3.1). The CRPropa framework allows modeling the transport behavior of energetic particles by either solving the Lorentz-Newton equation (so-called ballistic propagation) or by applying a set of stochastic differential equations (SDEs) to solve the GCR transport equation.
@@ 2 Modelling Framework
2.1 Magnetic field configuration
In the simulation, different magnetic field configurations can be used. The baseline model considered here is the Dipole + Quadrupole + Current Sheet (DQCS) model introduced by Banaszkiewicz_DQCS. This model consists of a dipole and quadrupole moment and a contribution from the current sheet in the Sun’s equatorial plane. It fits the solar magnetic field quite well in solar minimum conditions. A set of field lines is shown in Fig. 2. In CRPropa, the field is evaluated at each position based on the analytical description.
@@ 2.1 Magnetic field configuration
To evaluate the influence of the magnetic field configuration, we compare the results of the baseline field with the potential field source-surface (PFSS) model (Schatten69; Altschuler69) evaluated at different times. We calculate the PFSS field using the pfss Python package from Antony Yeates [^1] https://github.com/antyeates1983/pfss on a spherical grid containing 60 bins in $\log(r)$, 180 bins in $\cos\theta$, and 360 bins in $\phi$. The outer surface, at which the magnetic field is assumed to be purely radial, is set to $R_{ss}=2.5R_{\odot}$. In between the grid points the magnetic field is evaluated by tri-linear interpolation in spherical coordinates, although it is known that the interpolation might affect the results (Schlegel20).
The input of the PFSS field calculation is a magnetic map reconstructed from a 27-day solar rotation (Carrington Rotation, CR), i.e. synoptic or diachronic magnetic map. We use the synoptic maps measured with the Helioseismic and Magnetic Imager (HMI) (Schou-etal-2012) for the CR 2154, 2157, and 2224. The first two CR describe the active phase of the Sun, while the last one is close to the solar minimum. An illustration of the field line configuration in the different CRs is shown in Fig. 2.
2.2 Hadronic Interaction
@@
In this CRPropa module, it is assumed that all secondary particles are created in the propagation direction of the primary GCR. This is based on the assumption of ultra-relativistic particles within CRPropa. Griffith-etal-2025 introduce the effect of the opening angle in the interaction, which can be relevant for proton energies $E_{p}\lesssim 10\,\mathrm{GeV}$. In this work, we focus on gamma rays with $E_{\gamma}\gtrsim 1\,\mathrm{GeV}$, which are predominantly created by protons with $E_{p}\gtrsim 10\,\mathrm{GeV}$. Therefore, we do not expect a strong impact of this assumption.
Besides the cross-section, the other input needed is the target density. We use the chromosphere model from Avrett08. To extend the density profile to the source surface at $r=2.5\,R_{\odot}$, we follow the radial slope from Kontar19, but apply a lower normalization. The radial dependence of both models is shown in Fig. 3 as a function of the height above the solar surface.
@@ 2.3 CRPropa setup
Figure 4: Parametrisation of the local interstellar spectrum with the data measured by AMS-02 (AMS), CALET (Calet), CREAM (Cream), Pamela (Pamela), Voyager (Voyager13; Voyager16; Voyager19) and DAMPE (Dampe).
2.4 Filtering of gamma-ray directions
The production of gamma rays is dominantly in the direction of the parent proton due to the high Lorentz boost. As many of the incoming GCR protons are travelling towards the Sun, most gamma rays are absorbed by the solar disk and can not be observed from the Earth. To account for this effect in our simulation, we perform a filtering of the produced gamma rays in the post-processing. Here, we differentiate between three modes illustrated in Fig. 1. For further analysis, we construct three datasets, where every selection includes all particles selected before:
@@ 2.4 Filtering of gamma-ray directions
    In the second selection (passing), we add all gamma rays with an ingoing momentum that will not hit the solar surface (labeled passing in Fig. 1). Those gamma rays are produced by GCR protons with a nearly perpendicular pitch-angle, but the point of interaction is far enough away from the solar surface. Here, we note that the number of particles added in this selection is rather small, as the density drops significantly for larger heights over the surface.
@@
3 Mono Energetic simulation
As a first step, we analyze simulations with a mono-energetic injection to investigate the impact of the energy on several observables, starting with the production height above the surface (Sec. 3.1), the initial pitch-angle distribution leading to gamma-ray production (Sec. 3.2)
3.1 Production height of gamma rays

### v2

@@ Abstract
The Sun produces a steady signal of high-energy gamma rays through interactions of Galactic cosmic rays (GCRs) with its atmosphere. Observations with Fermi-LAT and HAWC have revealed a gamma-ray flux significantly higher than early theoretical predictions, with unexpected temporal and spectral features that suggest a crucial role of the solar magnetic field. In this work, we model GCR-induced gamma-ray emission at the solar disk using the CRPropa framework with realistic hadronic interactions, chromospheric density profiles, and several magnetic field configurations over the solar cycle. This allows us to quantify the gamma-ray emission of the entire solar disk for different phases of the solar activity cycle and we present, for the first time, maps of the production locations of gamma rays on the solar surface. We consider both mono-energetic and realistic power-law injection spectra in a simplified dipole-quadrupole-current-sheet model and potential-field source surface (PFSS) extrapolations for Carrington rotations during solar maximum and minimum. Our results show that magnetic mirroring and large-scale field topology strongly affect the spectral shape and spatial distribution of the emission, with slightly enhanced fluxes predicted at solar minimum. While our simulated baseline fluxes remain below observations, additional effects, such as heavier nuclei, Parker-field mirroring, and deeper atmospheric interactions, could result in further enhancements of fluxes closer to observational values. Hadronic interactions do not only produce gamma rays but also neutrinos. We estimate the expected neutrino flux from the Sun based on our predictions. We find that the expected flux is slightly below current upper limits from IceCube.
@@ I Introduction
The dominant high-energy cosmic ray component in the heliosphere is of Galactic origin. These Galactic Cosmic Rays (GCRs) are not only useful messengers from distant astrophysical systems (e.g., Becker Tjus and Merten, 2020), but they can also be used to infer information about the large-scale structure of the heliosphere (e.g., Pogorelov et al., 2017) as well as about the turbulence in the solar wind (e.g., Oughton and Engelbrecht, 2021; Fraternale et al., 2022).
@@
Even in the innermost heliosphere, high-energetic GCRs are important agents regarding the solar magnetic field. The deflection of GCRs in the latter produces the so-called solar cosmic-ray shadow (Becker Tjus et al., 2020; Aartsen et al., 2021, and references therein), which varies with solar activity and can, therefore, be used to infer information about the solar coronal magnetic field (e.g., Alfaro et al., 2024; Ng et al., 2025). The GCRs also interact with the atmosphere and the interior of the Sun. Seckel et al. (1991) were the first to estimate the fluxes of gamma-rays, neutrinos, antiprotons, neutrons, and antineutrons resulting from cascades that are initiated by the interactions of GCRs with the solar atmosphere. Aside from protons and heavier ions, GCR electrons and positrons can also produce gamma-rays via inverse Compton scattering of solar photons, as well as extreme ultraviolet (EUV) to hard X-ray emission via synchrotron radiation in the solar magnetic field. The transport of GCR electrons from 1 AU to the Sun and their subsequent synchrotron emission were recently modeled (Petrosian et al., 2023; Orlando et al., 2023). The steady flux of gamma-rays produced near the Sun as a result of GCR proton interactions is the primary interest in the present study.
@@
Gamma rays produced by the Sun in the GeV range are also associated with transient solar flare events. The maximum energy of flare-produced gamma rays is about 4 GeV (e.g. Ackermann et al., 2014), which is lower than most of the energy range discussed in the following, reaching well into the TeV regime. Regarding the steady gamma-ray flux of the Sun, one distinguishes two components of the GCR-induced gamma-ray emission, namely the disk and the halo component. The first one comes from the solar disk and was theoretically studied by Seckel et al. (1991) and detected first in EGRET data (Orlando and Strong, 2008). It results predominantly from GCR protons, which follow magnetic flux tubes in the solar atmosphere. When these particles interact with the solar atmosphere, they produce, via cascades, gamma-rays from the direction of the solar disk. The second, so-called halo component, results from inverse Compton scattering of cosmic-ray electrons in the extended solar atmosphere, quantified and validated by Orlando and Strong (2008, see also the review in ()) as far as possible with the EGRET data available at that time (see also Moskalenko et al., 2006). The predictions for the halo component and the corresponding EGRET observations have been confirmed with measurements made with the Large Area Telescope of the Fermi mission (Fermi-LAT, Abdo et al., 2011). The disk component was detected with Fermi as well, however, at a flux level about a factor of six higher (Nisa et al., 2019) than predicted by Seckel et al. (1991). These findings have been confirmed at lower energies in Ng et al. (2016) by extending the analysis of Fermi-LAT data from 1–10 GeV (Abdo et al., 2011) to the interval 1–100 GeV. With the High Altitude Water Cherenkov (HAWC) Observatory (DeYoung and HAWC Collaboration, 2012), solar disk gamma rays with higher energies in a low TeV range have been observed (Albert et al., 2023).
@@
Ng et al. (2016) identified a time-dependence of the flux of the disk component by a factor of two to three in the energy interval 1101-10 GeV, confirmed by Linden et al. (2022). This variation is present up to about 30 MeV, anti-correlated to solar activity, and cannot be explained exclusively with the corresponding modulation of GCRs, which amounts to $\sim$15% for 10 GeV protons and less for higher energies. So, as for the cosmic-ray shadow, the solar magnetic field must be expected to play a significant role: The field structure determines the trajectories of the GCRs in the solar atmosphere. First, it determines which particles initiate cascades that eventually lead to the formation of the disk component of gamma rays. And, second, the field also determines the temporal variation of the gamma-ray signal. Such variation with solar activity is also present in the HAWC data (Albert et al., 2023).
@@
Furthermore, Tang et al. (2018) reported an unexpected dip in the solar gamma-ray spectrum between 305030-50 GeV, which has recently been attributed to an energy cutoff in the turbulence-induced migration of GCRs into closed magnetic arcades (Puzzoni et al., 2024). While the confirmation of this needs further measurements, another finding remains, so far, unexplained. Linden et al. (2018) reported that the gamma-ray emission is not point-like and isotropic (Nisa et al., 2019) but exhibits a nearly constant polar component up to 100 GeV and an equatorial component that is brightest during solar minimum with energies above 200 GeV but not in the TeV-regime (Albert et al., 2018; Bartoli et al., 2019). Finally, the gamma-ray flux is characterized by power laws in energy, namely $E^{-2.2}$ below 200 GeV (Linden et al., 2022, and references therein), steepening to $E^{-3.6}$ above 500 GeV (Albert et al., 2023).
@@
Following the original idea by Seckel et al. (1991), recently Li et al. (2024a) have attempted to identify the magnetic field structures at the solar surface that are crucial for reflecting hadronic GCRs and have modeled the corresponding gamma-ray emission. To this end, the authors considered a solar surface magnetic field consisting of vertical flux tubes, representing those that form the so-called network elements, and of vertical flux sheets, representing those at the so-called down-flow lanes between granules. This way, the discrepancy between simulations and observations could be somewhat reduced, but it remains at a factor of two to four. While this model is quite advanced on a small scale, it assumes a spherically symmetric solar atmosphere on a large scale and, thus, does not take into account the large-scale coronal and interplanetary magnetic field structure. In the context of gamma-ray production, the coronal field very close to the Sun was taken into account by Hudson et al. (2020) using an MHD model and by Li et al. (2024b) employing the potential field source surface (PFSS) approach, again resulting in too low simulated fluxes.
@@
All of the above models did not consider a small-scale turbulent component of the large-scale solar magnetic field. This was done recently by Puzzoni et al. (2024), focusing on closed turbulent magnetic field arcades in the low photosphere on the gamma ray flux. This study was complemented by Puzzoni et al. (2025) with modeling a synthetic large-scale magnetic field with static, laminar open field lines in the chromosphere that are increasingly braided near the solar surface. A key finding of these studies is that the tension between the modelled and the observed gamma ray flux from the solar disk reduces with increasing turbulence level. In extension of their earlier work, Li et al. (2025) also included the effect of turbulence. Interestingly, the Alfvénic turbulence considered therein was found to reduce rather than increase the modelled flux.
@@
Extending the scale to a few solar radii, Hutchinson et al. (2022) considered the interplanetary Parker field and the turbulence-induced scattering of GCRs within, however, only in the context of solar flares and not yet computing gamma ray fluxes.
@@
II Modelling Framework
We model the gamma-ray emission with CRPropa 3.2 (Alves Batista et al., 2022), a publicly available tool for the propagation and interaction of charged particles in magnetic fields. While the code was originally written for ultra high energy cosmic rays (UHECRs) (Alves Batista et al., 2016), an extension for lower energies to cover GCRs has been introduced later on (Merten et al., 2017). The CRPropa framework allows modeling the transport behavior of energetic particles by either solving the Lorentz-Newton equation (so-called ballistic propagation) or by applying a set of stochastic differential equations (SDEs) to solve the GCR transport equation.
@@ II Modelling Framework
II.1 Magnetic field configuration
In the simulation, different magnetic field configurations can be used. The baseline model considered here is the Dipole + Quadrupole + Current Sheet (DQCS) model introduced by Banaszkiewicz et al. (1998). This model consists of a dipole and quadrupole moment and a contribution from the current sheet in the Sun’s equatorial plane. It fits the solar magnetic field quite well in solar minimum conditions. A set of field lines is shown in Fig. 2. In CRPropa, the field is evaluated at each position based on the analytical description.
@@ II.1 Magnetic field configuration
To evaluate the influence of the magnetic field configuration, we compare the results of the baseline field with the potential field source-surface (PFSS) model (Schatten et al., 1969; Altschuler and Newkirk, 1969) evaluated at different times. We calculate the PFSS field using the pfss Python package from Antony Yeates [^1] https://github.com/antyeates1983/pfss on a spherical grid containing 60 bins in $\log(r)$, 180 bins in $\cos\theta$, and 360 bins in $\phi$. The outer surface, at which the magnetic field is assumed to be purely radial, is set to $R_{ss}=2.5R_{\odot}$. In between the grid points the magnetic field is evaluated by tri-linear interpolation in spherical coordinates, although it is known that the interpolation might affect the results (Schlegel et al., 2020).
The input of the PFSS field calculation is a magnetic map reconstructed from a 27-day solar rotation (Carrington Rotation, CR), i.e. synoptic or diachronic magnetic map. We use the synoptic maps measured with the Helioseismic and Magnetic Imager (HMI) (Schou et al., 2012) for the CR 2154, 2157, and 2224. The first two CR describe the active phase of the Sun, while the last one is close to the solar minimum. An illustration of the field line configuration in the different CRs is shown in Fig. 2.
II.2 Hadronic Interaction
@@
In this CRPropa module, it is assumed that all secondary particles are created in the propagation direction of the primary GCR. This is based on the assumption of ultra-relativistic particles within CRPropa. Griffith et al. (2025) introduce the effect of the opening angle in the interaction, which can be relevant for proton energies $E_{p}\lesssim 10\,\mathrm{GeV}$. In this work, we focus on gamma rays with $E_{\gamma}\gtrsim 1\,\mathrm{GeV}$, which are predominantly created by protons with $E_{p}\gtrsim 10\,\mathrm{GeV}$. Therefore, we do not expect a strong impact of this assumption.
Besides the cross-section, the other input needed is the target density. We use the chromosphere model from Avrett and Loeser (2008). To extend the density profile to the source surface at $r=2.5\,R_{\odot}$, we follow the radial slope from Kontar et al. (2019), but apply a lower normalization. The radial dependence of both models is shown in Fig. 3 as a function of the height above the solar surface.
@@ II.3 CRPropa setup
Figure 4: Parametrisation of the local interstellar spectrum with the data measured by AMS-02 (Aguilar et al., 2016), CALET (Adriani et al., 2022), CREAM (Choi et al., 2022), Pamela (Adriani et al., 2013), Voyager (Stone et al., 2013; Cummings et al., 2016; Stone et al., 2019) and DAMPE (An et al., 2019).
II.4 Filtering of gamma-ray directions
The production of gamma rays is dominantly in the direction of the parent proton due to the high Lorentz boost. As many of the incoming GCR protons are travelling towards the Sun, most gamma rays are absorbed by the solar disk and can not be observed from the Earth. The absorption of gamma-rays is not taken into account within the CRPropa simulation, but can be treated in the post-processing. To account for this effect in our simulation, we perform a filtering of the produced gamma rays in the post-processing. Here, we differentiate between three modes illustrated in Fig. 1. For further analysis, we construct three datasets, where every selection includes all particles selected before:
@@ II.4 Filtering of gamma-ray directions
    In the second selection (passing), we add all gamma rays with an ingoing momentum that will not hit the solar surface (labeled passing in Fig. 1). Those gamma rays are produced by GCR protons with a nearly perpendicular pitch-angle, but the point of interaction is far enough away from the solar surface. Here, we note that the number of particles added in this selection is rather small, as the density drops significantly for larger heights over the surface. Nearly all results shown later on, are the same for the outgoing and passing selections. We refer the reader to focus on the passing results.
@@
III Mono Energetic simulation
As a first step, we analyze simulations with a mono-energetic injection to investigate the impact of the energy on several observables, starting with the production height above the surface (Sec. III.1), the trapping time of GCRs (Sec. III.2), the initial pitch-angle distribution leading to gamma-ray production (Sec. III.3) and their spatial distribution on the solar surface (Sec. III.4).
III.1 Production height of gamma rays